A charge of q = -5.3 nC is accelerated from rest through a potential


difference of 25 V. What is the change in electric potential energy of


the charge?

1
Expert's answer
2022-03-10T11:16:12-0500

The change in electric potential energy of the charge

"\Delta E_p=qV\\
=-5.3*10^{-9}*25=-1.3*10^{-7}\:\rm J"